OpenStudy (anonymous):

In a right triangle, the hypotenuse is 17 inches and a leg is 15 inches. Find the length of the third leg.

OpenStudy (anonymous):

Multiply 17^2 and 15^2. subtract 15^2 from 17^2. Next find the square root. Did you get an answer of 8?(:

OpenStudy (anonymous):

17^2 - 15^2 = b^2 b = 64^2 = 4096 4096 should be your answer. It's the pythaogrean theorem correct? So a^2 + b^2 = c^2 Vice versa OR c^2 - b^2 = a^2 c^2 - a^2 = b^2 Notice that if you have the hypotenuse which is C, then you always subtract either the a^2 or the b^2 to find the other value.

OpenStudy (anonymous):

Fang02, I believe you are wrong because if you check your answer 15^2+8^2=17^2. 225+64+289.